Events A & B are independent. Suppose Event A occurs with a probability 0.60 and Event B occurs with a probability of 0.74.
A) Compute the probability that B occurs but A does not occur.
B) Compute the probability that either A occurs without B or A and B both occur.

Answer by ikleyn(52798)

.
Events A & B are independent. Suppose Event A occurs with a probability 0.60 and
Event B occurs with a probability of 0.74.
A) Compute the probability that B occurs but A does not occur.
B) Compute the probability that either A occurs without B or A and B both occur.
~~~~~~~~~~~~~~~~~~~

Since events A and B are independent, the probability

    P(A & B) = P(A)*(P(B) = 0.60*0.74 = 0.444.


(A)  the probability that B occurs but A does not occur = P(B) - P(A & B) = 0.74 - 0.444 = 0.296.    ANSWER


(B)  the probability that either A occurs without B or (A & B) both occur = 

         =                  (P(A) - P(A & B))       +   P(A & B)          = P(A) = 0.60.    ANSWER

Solved.